The length of the garden is 4 meters and 56 meters of fencing is required to completely enclose it. What is the width of the garden?
How do I do this?

Answer by caseyec(1)

if the garden is 4 meters long and 'x' represents the width the equation would be 56=4*2+x*2
4*2=8 56-8=48 48/2=24 The width of the garden is 24 meters